Sub 24
Ques. 16

Find the length of a side of each of two squares given that the sum of their perimeters is 44 ft and the sum of their areas is 73 ft^2.
1
Expert's answer
2012-05-15T11:22:04-0400
x^2+y^2 = 73
4x+4y = 44, x +y = 11, x = 11- y

(11 - y)^2 +y^2 = 73,
y^2 -11y +24 = 0, y1 = 3, y2 = 8
x1 = 8, x2 = 3.
So the sides are 3 ft and
8 ft